OpenStudy (anonymous):

Please help... A card is drawn from a deck of 10 cards numbered 1-10. The card is replaced in the deck and another card is drawn. Find the probability of each outcome.... P(5 and then a 3) P(two even numbers) P(two numbers greater than 4) P(6 and then an odd number) Please help

OpenStudy (anonymous):

Also, P(an odd number and then an even number, P(a number greater than 7 and then a number less than 6) What is the probability of tossing a coin three times and getting heads each time (wouldn't it just be 1/2?)

OpenStudy (anonymous):

My class is reviewing this and I don't get it. At all....Even tho I learned it, it has been awhile since I have seen it.... :(

OpenStudy (anonymous):

Pretty sure the first one is a 1/100 probability second in 1/4 (The idea is pulling a 5 is a 1 in 10 chance pulling a 3 is a 1 in 10 so pulling one then the other you multiply 1/10 by 1/10 giving you 1/100) (same as the next pulling even is 5/10 simplified to 1/2. 1/2 times 1/2 is 1/4)

OpenStudy (anonymous):

Most of the answers in the back of the book is a #/20

OpenStudy (anonymous):

so following those same rules there are 6 cards greater then 10 so (6/10)*(6/10) which is 9/25 or 36% and the 4th one is (1/10)*(1/2)

OpenStudy (anonymous):

You have to have the answer in the "language" of fractions.

OpenStudy (anonymous):

But we have to show our work also.

OpenStudy (anonymous):

ok so are any of these checkable? because you can just divide the 1/100 by 5 is .2/20 and the 1/4 you times by 5 and would get 5/20 but im not sure exactly what your teacher is looking for

OpenStudy (anonymous):

give me a problem you have the answer to as well

OpenStudy (anonymous):

ok... one second.

OpenStudy (anonymous):

I don't know how the book got this but... P(6 and then an odd number)=1/20 I think that since there is 10 cards but they could be any like, um...like a diamond, club, spade, or heart....so it would be a total of 40 choices. I think....right?

OpenStudy (anonymous):

I'm not sure tho.

OpenStudy (anonymous):

And then if P(a number greater than 7 and then a number less than 6)=3/20, I don't know how they got it. It is very confusing.

OpenStudy (anonymous):

no because it doesn't specify suits. how you get the 1/20 is following the same steps. take how many card out of ten will satisfy both so to get a 6 there is 1 out of 10 so 1/10 to get an odd there is a 5 out of 10 so 5/10 reduced to 1/2. (1/10)*(1/2)=1/20

OpenStudy (anonymous):

But I am pretty sure you got the 1/4 right. Yeah, ya did. Oh so like you would just multiply the 2 answers you got and then you have the final answer???

OpenStudy (anonymous):

If so, I get it now.

OpenStudy (anonymous):

greater then 7 is a 3/10 chance less then 6 is a 5/10 chance multiply together and its a 15/100 divide by 5 and you get 3/20

OpenStudy (anonymous):

ya exactly

OpenStudy (anonymous):

ya you just take each probability and multiply them and it gives you final

OpenStudy (anonymous):

ok. thanks for the help. I really appreciate it.

OpenStudy (anonymous):

oh and the coin is (1/2)*(1/2)*1/2) so 1/8

OpenStudy (anonymous):

k. thanks!!!